# Zero-shot Hierarchical Plant Segmentation via Foundation Segmentation Models and Text-to-image Attention

### v2

@@ Abstract
Foundation segmentation models achieve reasonable leaf instance extraction from top-view crop images without training (i.e., zero-shot). However, segmenting entire plant individuals with each consisting of multiple overlapping leaves remains challenging. This problem is referred to as a hierarchical segmentation task, typically requiring annotated training datasets, which are often species-specific and require notable human labor. To address this, we introduce ZeroPlantSeg, a zero-shot segmentation for rosette-shaped plant individuals from top-view images. We integrate a foundation segmentation model, extracting leaf instances, and a vision-language model, reasoning about plants’ structures to extract plant individuals without additional training. Evaluations on datasets with multiple plant species, growth stages, and shooting environments demonstrate that our method surpasses existing zero-shot methods and achieves better cross-domain performance than supervised methods. Implementations are available at https://github.com/JunhaoXing/ZeroPlantSeg.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
(a) Ground truth mask. (b) SAM: $IoU=0.298$. (c) ZeroPlantSeg (ours): $IoU=\bm{0.883}$.
Figure 1: Zero-shot plant individual segmentation by Segment Anything Model (SAM) [13] and our ZeroPlantSeg. SAM inputs point prompts by randomly selecting $10$ points from the ground-truth plant individual mask, as indicated by the green stars in the figure.
@@ 1 Introduction
Computer vision is essential in image-based plant phenotyping, which extracts plants’ traits from images for applications in plant sciences and agriculture. In particular, instance segmentation of leaves and plants is actively utilized for capturing important phenotypic attributes such as plant count, area, and morphology. To this end, the joint segmentation of leaves and plant individuals is studied as hierarchical segmentation [30, 29, 24], where most existing methods rely on annotated training datasets, which are often species- or even environment-specific and thus limit the practical usefulness due to the required human labor.
Zero-shot segmentation models, such as the Segment Anything Model (SAM) [13], offer a promising alternative by eliminating the need for domain-specific training. In agricultural applications, SAM and its variants have been actively utilized to segment leaf instances (e.g., Leaf-only SAM [31]), where several heuristics of plant leaves are introduced for accurate segmentation. However, a straightforward application of zero-shot segmentation models is not reasonable for segmenting plant individuals, where their leaves have complex overlaps and mutual occlusions between individuals, as shown in Fig. 1(b).
To address these issues, we propose ZeroPlantSeg, a zero-shot solution for hierarchical plant segmentation, targeting the top-view images capturing rosette-shaped plants, which is a common assumption in existing methods [30, 29, 24] and datasets [28, 10, 2]. Similar to the existing zero-shot leaf segmentation methods (e.g., [31]), we use SAM and introduce plants’ heuristics to extract leaf instances in the scene. To segment the plant individuals, we first leverage a prior from vision language models (VLMs), where we extract the stem location from the leaf instances. Using the stem locations, we then perform a clustering-based method for combining the leaves belonging to the same individual.
Experiments on three datasets (Phenobench [28], GrowliFlower [10], and SB20 [2]) show that our method outperforms a straightforward application of zero-shot segmentation. Besides, our method shows a notably better capability on cross-domain settings (i.e., different species and environments) compared to existing supervised methods for hierarchical segmentation [30, 29, 24].
@@
  • We introduce the first zero-shot method for hierarchical plant segmentation, ZeroPlantSeg. Our pipeline addresses the challenge of segmenting overlapping plant structures without supervision by integrating pre-trained SAM for leaf segmentation and VLM-based attention for plant structure reasoning.
@@ 2 Related Work
Accurate segmentation of leaves and plants enables early identification of disease symptoms, such as spots, discoloration, or lesions on leaves. Also, it allows for precise measurement of leaf area, shape, and count, aiding in plant growth monitoring. However, classical segmentation methods relying on color-based [1] or brightness-based [19] thresholding struggle with the dynamic nature of plant growth, including variations in color, size, and structure. To overcome these limitations, learning-based approaches [7, 25] have been widely adopted, offering robustness against variations in lighting and plant morphology.
@@
Hierarchical segmentation, i.e., simultaneous plant and leaf level segmentation, has recently garnered interest due to its ability to provide granular information, which indicates growth stages [28]. Notable works [30, 29, 24] known as supervised methods have explored hierarchical segmentation by leveraging plant structures rather than independently training separate models for leaves and plants. Furthermore, a self-supervised method requiring fewer annotations [27] is proposed. However, these existing methods still require extensive annotated training data, incurring significant labor costs. Moreover, the generalizability of these models across plant species and diverse imaging conditions remains uncertain, as there are only limited datasets [20, 10, 28] for hierarchical segmentation.
@@
Zero-shot segmentation models offer a promising alternative by eliminating the need for domain-specific training [33, 15]. Those models trained on vast datasets encompassing diverse scenarios with Vision Transformer (ViT) [3, 17] backbones can segment unseen images with remarkable generalization ability. For instance, SAM [13] is pre-trained on 11 million images containing over 1 billion masks (i.e., SA1B dataset [13]), enabling mask generation based on arbitrary object prompts. Furthermore, SAM provides the Automatic Mask Generator (AMG) that produces masks for everything in the image using a regular grid of point prompts. However, SAM’s output is class-agnostic, making plant segmentation challenging in cluttered images. Methods integrating VLMs have emerged to introduce class awareness into segmentation outputs by learning adaptation for Contrastive Language-Image Pre-training (CLIP) [21] or cross-attention [32] between text and image features [6, 23, 26].
In agricultural applications, SAM has been utilized to segment leaf instances from laboratory-controlled images and UAV-acquired field imagery. Leaf-only SAM proposed by Williams et al[31] uses AMG on an image of potato plants and uses color, area, and shape filters to choose only leaf masks from class-agnostic masks. Torres-Lomas et al[4] use AMG of SAM to segment grape berries from a bunch and measure their number and area. Yu et al[8] capture images of a soybean canopy using a UAV, remove the background using a machine learning-based method, and segment individual leaves with SAM to perform variety classification based on leaf morphology. However, its performance degrades when applied to complex outdoor images, often merging multiple plants into a single mask. Fine-tuning SAM with object detection models or employing additional prompt-based refinements has been explored, but these approaches still require training data.
@@ 3.1 Leaf Instance Segmentation
We use the Automatic Mask Generator (AMG) in SAM [13] to obtain the candidates of leaf masks from an input image. AMG generates masks from point prompts placed equally distributed on the image; however, obtaining masks for small instances compared to the image size is challenging, which is likely to occur in images captured from UAVs. We thus crop and upsample the input image using a sliding window to obtain masks for small leaves. Masks located at the window boundaries are not used.
SAM does not classify the masks it produces, and its output contains masks for all class objects in the image. Therefore, we extract leaf masks by OVSeg [14], a mask-adapted image classifier that takes the text and masked image as input. It encodes the mask and text prompt with a mask-adapted image encoder and text encoder, respectively, and calculates their cosine similarity to predict each mask’s class. We use the text prompts “green leaf” and “soil” expecting that the scene only contains plants and soil. Masks with a larger similarity with “soil” than with “green leaf” are discarded. As shown in Fig. 3(b), we obtain binary masks of leaf, and we describe the set of those masks as $\mathcal{M}=\{\bm{M}^{i}\}_{i=1}^{N}$, where $\bm{M}^{i}\in\{0,1\}^{H\times W}$.
@@
The masks by AMG have overlapping regions, i.e., masks for two or more leaves are produced. Besides, since we use a sliding window to slice an image before inputting it to SAM, we have duplicated masks of the same instance in the overlapping region of the windows. To obtain unified leaf instances, we remove these duplicated masks. We use NMS in Leaf Only SAM [31], which combines or suppresses the overlapping masks. For details, please refer to [31].
@@ 3.2 Plant Individual Segmentation
We then extract the stem lines in the leaf images $\hat{\mathcal{L}}$. We feed each $\hat{\bm{L}}^{i}$ to the text-to-image cross-attention module of the pre-trained VLM with a text prompt that describes the stem-like part of a plant, namely, “stem” and “petiole”. We use the output features of the text-to-image cross-attention (i.e., SwinT OGC) module of Grounding DINO [16]. The features are spanned in four resolutions from $8\times 8$ to $64\times 64$. We resize and average them to obtain the final feature map $\bm{F}^{i}\in\mathbb{R}^{100\times 100}_{+}$, as shown in Fig. 4(a).
@@
We use unsupervised clustering on the set of keypoints $\mathcal{Q}$ to obtain plant instances. Although the most naive way is to apply a simple clustering method such as DBSCAN [18], it can produce invalid clusters since $\mathcal{Q}$ includes both leaf tips and bases, where leaf tips are far from the plants’ center. To address this issue, we perform the clustering in a greedy manner, assuming many leaf base points are located close to the center of the plants’ individuals in the rosette-shaped plants.
@@ 4 Experiments
We compare our ZeroPlantSeg with supervised hierarchical segmentation by Weyler et al[29] and Gianmarco et al[24], which we hereafter call Weyler et al. and HAPT, respectively. Besides, we also assess zero-shot segmentation models that do not require training data like our method, including LeafOnlySAM [31], Grounded SAM [23], and Grounded SAM 2 [22].
@@ 4.3 Baselines
We manually train the two supervised models on the PhenoBench [28] dataset. We train HAPT [24] with AdamW [9] for $200$ epochs with a batch size of $8$. We set the step learning rate scheduler with an initial value of $4\times 10^{-4}$ for backbone, and three exponential schedulers with initial learning rates of ($4\times 10^{-4}$, $8\times 10^{-4}$, and $8\times 10^{-4}$) for the semantic, plant instance, and leaf instance decoder, respectively. During training, we apply dropout with a probability of $0.15$. We train Weyler et al. [29] with Adam [11] for $512$ epochs with a batch size of $1$. We set the initial learning rate to $1\times 10^{-3}$ and subsequently apply a polynomial learning rate decay $(1-\frac{e}{512})^{0.9}$, where $e$ is the current epoch.
@@
We use the zero-shot models, including Grounded SAM (G-SAM) [23], Grounded SAM 2 (G-SAM2) [22], and LeafOnlySAM (LO-SAM) [31] with the pre-trained weights provided in their official repositories. For the grounded version of SAMs (i.e., G-SAM [23] and G-SAM2 [22]), we use two text prompts, “leaf” and “plant”, to obtain leaf and plant individual masks, respectively. If a pixel is contained in multiple leaf masks, we assign the smallest one similarly to LO-SAM [31]. For plant individual masks, we assigned the largest mask for each pixel. For leaf segmentation, predicted masks that cover more than 30 % of the images are discarded. For plant segmentation, those that are larger than 50 % are discarded. LO-SAM [31] is literally designed for segmenting leaf only; thus, we do not assess the accuracy for the plant individual segmentation.
@@ 4.4 Metrics
We assess the standard metrics for segmentation accuracies: precision, recall, and average precision (AP). For all metrics, we use the Intersection over Union (IoU) threshold of $0.5$ and denote them as $Prec_{50}$$Rec_{50}$, and $AP_{50}$, respectively. Following previous research of hierarchical plant segmentation [28], we also evaluate the result with the Panoptic Quality (PQ) [12], which assesses both segmentation quality and recognition quality. We compute $PQ$ for plant instances and leaf instances, and denote them as $PQ_{\text{plant}}$ and $PQ_{\text{leaf}}$, respectively.
@@ 4.5 Results
We describe the experimental results for PhenoBench, GrowliFlower, and SB20 datasets. Since the supervised baselines (i.e., HAPT [24] and Weyler et al[29]) are trained on the PhenoBench dataset, testing on GrowliFlower contains the domain gap between the training and testing datasets on plant species, and testing on the SB20 dataset contains the domain gap on the shooting environment. For each dataset, HAPT and Grounded SAM are selected as representative ones for supervised models and zero-shot models, respectively. For comparisons with other models (i.e., G-SAM2 [22] and LO-SAM [31]), please refer to the supplementary materials.
@@
HAPT [24] $66.75$ $45.96$ $37.40$ $41.66$ $50.19$
Weyler et al[29] $44.26$ $35.71$ $24.63$ $32.12$ $43.13$
G-SAM [23] - $\bm{65.21}$ $14.89$ $13.80$ $19.95$ $11.12$
G-SAM2 [22] - $18.80$ $30.08$ $15.89$ $18.27$ $10.19$
LO-SAM [31] - $9.79$ $13.50$ $4.99$ - $7.42$
@@
Table 2: Evaluation of each day by $AP_{50}$ on the PhenoBench dataset [28].
@@
HAPT [24] $47.84$ $39.08$ $22.55$
Weyler et al[29] $31.84$ $30.89$ $14.65$
G-SAM [23] - $12.80$ $18.68$ $11.38$
G-SAM2 [22] - $14.12$ $15.36$ $20.98$
LO-SAM [31] - $2.74$ $5.39$ $7.44$
@@
Figure 5 shows the visual results on the PhenoBench dataset, and Tab. 1 shows the quantitative evaluation. Our method with predicted leaf masks outperforms zero-shot models without any training data required, and achieves comparable accuracies to a supervised method (i.e., Weyler et al[29]).
For further analysis, an evaluation of the results by $AP_{50}$ for each growth stage of the crop (date of dataset captured) is shown in Tab. 2. Overall, our method consistently achieves better accuracies than the unsupervised method (i.e., G-SAM [23]). Interestingly, our method achieves the best result on June 5, which is the most challenging case that contains complex overlaps among leaves (see Fig. 5).
@@
HAPT [24] - $6.45$ $3.93$ $14.60$ $15.70$ $2.57$
Weyler et al[29] - $0.06$ $0.12$ $0.10$ $0.03$ $0.00$
G-SAM [23] - $80.97$ $64.06$ 54.954.92 $48.90$ $26.65$
G-SAM2 [22] - $25.00$ $16.67$ $18.18$ $20.06$ $15.38$
LO-SAM [31] - $2.64$ $2.38$ $0.84$ - $1.47$
@@
HAPT [24] - $53.24$ $8.17$ $7.62$ $11.89$
Weyler et al[29] - $0.00$ $0.00$ $0.00$ $0.00$
G-SAM [23] - $\bm{86.56}$ $18.75$ $18.22$ $25.47$
G-SAM2 [22] - $43.58$ $29.29$ $22.47$ $24.79$

### v3

@@ Abstract
Foundation segmentation models achieve reasonable leaf instance extraction from top-view crop images without training (i.e., zero-shot). However, segmenting entire plant individuals with each consisting of multiple overlapping leaves remains challenging. This problem is referred to as a hierarchical segmentation task, typically requiring annotated training datasets that are often species-specific and require significant human labor. To address this, we introduce ZeroPlantSeg, a zero-shot segmentation for rosette-shaped plant individuals from top-view images. We integrate a foundation segmentation model, extracting leaf instances, and a vision-language model, reasoning about plants’ structures to extract plant individuals without additional training. Evaluations on datasets with multiple plant species, growth stages, and shooting environments demonstrate that our method surpasses existing zero-shot methods and achieves better cross-domain performance than supervised methods. Implementations are available at https://github.com/JunhaoXing/ZeroPlantSeg.
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
(a) Ground-truth mask (b) G-SAM [GSAM] (c) G-SAM 2 [ravi2024sam2segmentimages] (d) ZeroPlantSeg (Ours)
Figure 1: Zero-shot hierarchical plant segmentation by our ZeroPlantSeg in comparison with other zero-shot methods. Top: Leaf instance segmentation. Bottom: Plant individual segmentation. The IoU indicates the accuracy of plant individual segmentation. Our method achieves compelling results for both leaf instance and plant individual segmentation, validated by the high IoU score.
@@ 1 Introduction
Computer vision is essential in image-based plant phenotyping, which extracts plants’ traits from images for applications in plant sciences and agriculture. In particular, instance segmentation of leaves and plants is actively utilized for capturing important phenotypic attributes such as plant count, area, and morphology. To this end, the joint segmentation of leaves and plant individuals is studied as hierarchical segmentation [weylerJointPlantLeaf2022, weylerBaseline, HAPT]. Most existing methods rely on annotated training datasets, which are often species- or environment-specific. Thus, their practical usefulness is limited due to the required human labor.
Zero-shot segmentation models, such as the Segment Anything Model (SAM) [SAM], offer a promising alternative by eliminating the need for domain-specific training. In agricultural applications, SAM and its variants have been actively utilized to segment leaf instances (e.g., Leaf-only SAM [LSAM]), where several heuristics of plant leaves are introduced for accurate segmentation. However, a straightforward application of zero-shot segmentation models is not reasonable for segmenting plant individuals, where their leaves have complex overlaps and mutual occlusions between individuals, as shown in Fig. 1(b).
To address these issues, we propose ZeroPlantSeg, a zero-shot solution for hierarchical plant segmentation, targeting the top-view images capturing rosette-shaped plants, which is a common assumption in existing methods [weylerJointPlantLeaf2022, weylerBaseline, HAPT] and datasets [PhenoBench, Growliflower, SB20]. Similar to the existing zero-shot leaf segmentation methods (e.g., [LSAM]), we use SAM and introduce plants’ heuristics to extract leaf instances in the scene. To segment the plant individuals, we first leverage a prior from vision language models (VLMs), where we extract the stem location from the leaf instances. Using the stem locations, we then perform a clustering-based method for combining the leaves belonging to the same individual.
Experiments on three datasets (Phenobench [PhenoBench], GrowliFlower [Growliflower], and SB20 [SB20]) show that our method outperforms a straightforward application of zero-shot segmentation. Besides, our method shows a notably better capability on cross-domain settings (i.e., different species and environments) compared to existing supervised methods for hierarchical segmentation [weylerJointPlantLeaf2022, weylerBaseline, HAPT].
@@
  • We introduce the first zero-shot method for hierarchical plant segmentation, ZeroPlantSeg. Our pipeline addresses the challenge of segmenting overlapping plant structures without supervision by integrating a pre-trained SAM for leaf segmentation and a VLM-based attention for plant structure reasoning.
@@ 2 Related Work
Accurate segmentation of leaves and plants enables early identification of disease symptoms, such as spots, discoloration, or lesions on leaves. Also, it allows for precise measurement of leaf area, shape, and count, aiding in plant growth monitoring. However, classical segmentation methods relying on color-based [achanta2010slic] or brightness-based [Meyer1992ColorIS] thresholding struggle with the dynamic nature of plant growth, including variations in color, size, and structure. To overcome these limitations, learning-based approaches [he2017mask, romera2016recurrent] have been widely adopted, offering robustness against variations in lighting and plant morphology.
@@
Hierarchical segmentation, i.e., simultaneous plant and leaf level segmentation, has recently garnered interest due to its ability to provide granular information, which indicates growth stages [PhenoBench]. Notable works [weylerJointPlantLeaf2022, weylerBaseline, HAPT] known as supervised methods have explored hierarchical segmentation by leveraging plant structures rather than independently training separate models for leaves and plants. Furthermore, a self-supervised method requiring fewer annotations [weylerPanopticSegmentationPartial2024] is proposed. However, these existing methods still require extensive annotated training data, incurring significant labor costs. Moreover, the generalizability of these models across plant species and diverse imaging conditions remains uncertain, as there are only limited datasets [CVPPPdataset, Growliflower, PhenoBench] for hierarchical segmentation.
@@
Zero-shot segmentation models offer a promising alternative by eliminating the need for domain-specific training [SEEM, Segnext]. Those models trained on vast datasets encompassing diverse scenarios with Vision Transformer (ViT) [ViT, SwinT] backbones can segment unseen images with remarkable generalization ability. For instance, SAM [SAM] is pre-trained on 11 million images containing over 1 billion masks (i.e., SA1B dataset [SAM]), enabling mask generation based on arbitrary object prompts. Furthermore, SAM provides the Automatic Mask Generator (AMG) that produces masks for everything in the image using a regular grid of point prompts. However, SAM’s output is class-agnostic, making plant segmentation challenging in cluttered images. Methods integrating VLMs have emerged to introduce class awareness into segmentation outputs by learning adaptation for Contrastive Language-Image Pre-training (CLIP) [CLIP] or cross-attention [xu2015crossatt] between text and image features [OVSAM, GSAM, HIPIE].
In agricultural applications, SAM has been utilized to segment leaf instances from laboratory-controlled images and UAV-acquired field imagery. Leaf-only SAM proposed by Williams et al. [LSAM] uses AMG on an image of potato plants and uses color, area, and shape filters to choose only leaf masks from class-agnostic masks. Torres-Lomas et al. [TorresLomasSamForComprehensive] use AMG of SAM to segment grape berries from a bunch and measure their number and area. Yu et al. [YuTimeSeriesField] capture images of a soybean canopy using a UAV, remove the background using a machine learning-based method, and segment individual leaves with SAM to perform variety classification based on leaf morphology. However, its performance degrades when applied to complex outdoor images, often merging multiple plants into a single mask. Fine-tuning SAM with object detection models or employing additional prompt-based refinements has been explored, but these approaches still require training data.
@@ 3.1 Leaf Instance Segmentation
We use the Automatic Mask Generator (AMG) in SAM [SAM] to obtain the candidates of leaf masks from an input image. AMG generates masks from point prompts placed equally distributed on the image; however, obtaining masks for small instances compared to the image size is challenging, which is likely to occur in images captured from UAVs. We thus crop and upsample the input image using a sliding window to obtain masks for small leaves. Masks located at the window boundaries are not used.
SAM does not classify the masks it produces, and its output contains masks for all class objects in the image. Therefore, we extract leaf masks by OVSeg [OVSeg], a mask-adapted image classifier that takes the text and masked image as input. It encodes the mask and text prompt with a mask-adapted image encoder and text encoder, respectively, and calculates their cosine similarity to predict each mask’s class. We use the text prompts “green leaf” and “soil” expecting that the scene only contains plants and soil. Masks with a larger similarity with “soil” than with “green leaf” are discarded. As shown in Fig. 3(b), we obtain binary masks of leaf, and we describe the set of those masks as $\mathcal{M}=\{\bm{M}^{i}\}_{i=1}^{N}$, where $\bm{M}^{i}\in\{0,1\}^{H\times W}$.
@@
The masks by AMG have overlapping regions, i.e., masks for two or more leaves are produced. Besides, since we use a sliding window to slice an image before inputting it to SAM, we have duplicated masks of the same instance in the overlapping region of the windows. To obtain unified leaf instances, we remove these duplicated masks. We use NMS in Leaf Only SAM [LSAM], which combines or suppresses the overlapping masks. For details, please refer to [LSAM].
@@ 3.2 Plant Individual Segmentation
We then extract the stem lines in the leaf images $\hat{\mathcal{L}}$. We feed each $\hat{\bm{L}}^{i}$ to the text-to-image cross-attention module of the pre-trained VLM with a text prompt that describes the stem-like part of a plant, namely, “stem” and “petiole”. We use the output features of the text-to-image cross-attention (i.e., SwinT OGC) module of Grounding DINO [GDINO]. The features are spanned in four resolutions from $8\times 8$ to $64\times 64$. We resize and average them to obtain the final feature map $\bm{F}^{i}\in\mathbb{R}^{100\times 100}_{+}$, as shown in Fig. 4(a).
@@
We use unsupervised clustering on the set of keypoints $\mathcal{Q}$ to obtain plant instances. Although the most naive way is to apply a simple clustering method such as DBSCAN [DBSCAN], it can produce invalid clusters since $\mathcal{Q}$ includes both leaf tips and bases, where leaf tips are far from the plants’ center. To address this issue, we perform the clustering in a greedy manner, assuming many leaf base points are located close to the center of the plants’ individuals in the rosette-shaped plants.
@@ 4 Experiments
We compare our ZeroPlantSeg with supervised hierarchical segmentation by Weyler et al. [weylerBaseline] and Gianmarco et al. [HAPT], which we hereafter call Weyler et al. and HAPT, respectively. Besides, we also assess zero-shot segmentation models that do not require training data like our method, including LeafOnlySAM [LSAM], Grounded SAM [GSAM], and Grounded SAM 2 [ravi2024sam2segmentimages].
@@ 4.3 Baselines
We manually train the two supervised models on the PhenoBench [PhenoBench] dataset. We train HAPT [HAPT] with AdamW [AdamW] for $200$ epochs with a batch size of $8$. We set the step learning rate scheduler with an initial value of $4\times 10^{-4}$ for backbone, and three exponential schedulers with initial learning rates of ($4\times 10^{-4}$, $8\times 10^{-4}$, and $8\times 10^{-4}$) for the semantic, plant instance, and leaf instance decoder, respectively. During training, we apply dropout with a probability of $0.15$. We train Weyler et al. [weylerBaseline] with Adam [Adam] for $512$ epochs with a batch size of $1$. We set the initial learning rate to $1\times 10^{-3}$ and subsequently apply a polynomial learning rate decay $(1-\frac{e}{512})^{0.9}$, where $e$ is the current epoch.
@@
We use the zero-shot models, including Grounded SAM (G-SAM) [GSAM], Grounded SAM 2 (G-SAM2) [ravi2024sam2segmentimages], and LeafOnlySAM (LO-SAM) [LSAM] with the pre-trained weights provided in their official repositories. For the grounded version of SAMs (i.e., G-SAM [GSAM] and G-SAM2 [ravi2024sam2segmentimages]), we use two text prompts, “leaf” and “plant”, to obtain leaf and plant individual masks, respectively. If a pixel is contained in multiple leaf masks, we assign the smallest one similarly to LO-SAM [LSAM]. For plant individual masks, we assigned the largest mask for each pixel. For leaf segmentation, predicted masks that cover more than 30 % of the images are discarded. For plant segmentation, those that are larger than 50 % are discarded. LO-SAM [LSAM] is literally designed for segmenting leaf only; thus, we do not assess the accuracy for the plant individual segmentation.
@@ 4.4 Metrics
We assess the standard metrics for segmentation accuracy: precision, recall, and average precision (AP). For all metrics, we use the Intersection over Union (IoU) threshold of $0.5$ and denote them as $Prec_{50}$$Rec_{50}$, and $AP_{50}$, respectively. Following previous research of hierarchical plant segmentation [PhenoBench], we also evaluate the result with the Panoptic Quality (PQ) [PanopticSegmentation], which assesses both segmentation quality and recognition quality. We compute $PQ$ for plant instances and leaf instances, and denote them as $PQ_{\text{plant}}$ and $PQ_{\text{leaf}}$, respectively.
@@ 4.5 Results
We describe the experimental results for PhenoBench, GrowliFlower, and SB20 datasets. Since the supervised baselines (i.e., HAPT [HAPT] and Weyler et al. [weylerBaseline]) are trained on the PhenoBench dataset, testing on GrowliFlower contains the domain gap between the training and testing datasets on plant species, and testing on the SB20 dataset contains the domain gap on the shooting environment. For each dataset, HAPT and Grounded SAM are selected as representative ones for supervised models and zero-shot models, respectively. For comparisons with other models (i.e., G-SAM2 [ravi2024sam2segmentimages] and LO-SAM [LSAM]), please refer to the supplementary materials.
@@
HAPT [HAPT] $66.75$ $45.96$ $37.40$ $41.66$ $50.19$
Weyler et al. [weylerBaseline] $44.26$ $35.71$ $24.63$ $32.12$ $43.13$
G-SAM [GSAM] - $\bm{65.21}$ $14.89$ $13.80$ $19.95$ $11.12$
G-SAM2 [ravi2024sam2segmentimages] - $18.80$ $30.08$ $15.89$ $18.27$ $10.19$
LO-SAM [LSAM] - $9.79$ $13.50$ $4.99$ - $7.42$
@@
Table 2: Evaluation of each day by $AP_{50}$ on the PhenoBench dataset [PhenoBench].
@@
HAPT [HAPT] $47.84$ $39.08$ $22.55$
Weyler et al. [weylerBaseline] $31.84$ $30.89$ $14.65$
G-SAM [GSAM] - $12.80$ $18.68$ $11.38$
G-SAM2 [ravi2024sam2segmentimages] - $14.12$ $15.36$ $20.98$
LO-SAM [LSAM] - $2.74$ $5.39$ $7.44$
@@
Figure 5 shows the visual results on the PhenoBench dataset, and Tab. 1 shows the quantitative evaluation. Our method with predicted leaf masks outperforms zero-shot models without any training data required, and achieves comparable accuracies to a supervised method (i.e., Weyler et al. [weylerBaseline]).
For further analysis, an evaluation of the results by $AP_{50}$ for each growth stage of the crop (date of dataset captured) is shown in Tab. 2. Overall, our method consistently achieves better accuracies than the unsupervised method (i.e., G-SAM [GSAM]). Interestingly, our method achieves the best result on June 5, which is the most challenging case that contains complex overlaps among leaves (see Fig. 5).
@@
HAPT [HAPT] - $6.45$ $3.93$ $14.60$ $15.70$ $2.57$
Weyler et al. [weylerBaseline] - $0.06$ $0.12$ $0.10$ $0.03$ $0.00$
G-SAM [GSAM] - $80.97$ $64.06$ 54.954.92 $48.90$ $26.65$
G-SAM2 [ravi2024sam2segmentimages] - $25.00$ $16.67$ $18.18$ $20.06$ $15.38$
LO-SAM [LSAM] - $2.64$ $2.38$ $0.84$ - $1.47$
@@
HAPT [HAPT] - $53.24$ $8.17$ $7.62$ $11.89$
Weyler et al. [weylerBaseline] - $0.00$ $0.00$ $0.00$ $0.00$
G-SAM [GSAM] - $\bm{86.56}$ $18.75$ $18.22$ $25.47$
G-SAM2 [ravi2024sam2segmentimages] - $43.58$ $29.29$ $22.47$ $24.79$